### v1

# Characterizing the Residual Finiteness of Conjugation Quandles of the Baumslag-Solitar Groups

Abstract.
This article fully characterize the residual finiteness of conjugation quandles of the Baumslag-Solitar groups. The key difference between groups and their derived conjugation quandles is that many conjugation quandles are infinitely generated even when their underlying groups are finitely generated. Notably, we prove that the conjugation quandles of Baumslag-Solitar groups are always infinitely generated. Moreover, we show that if a conjugation quandle is Hopfian, then its underlying group is also Hopfian. From this, we identity the conditions for non-Hopfian conjugation quandles of Baumslag-Solitar groups.
2020 Mathematics Subject Classification:
Primary 57K12, 20E26
1. Introduction
Quandles are algebraic structures based on Reidemeister moves in knot theory. Quandles were introduced independently by Joyce [11] and Matveev [13]. They have been applied in constructing invariants for knots (see, for example [7]) and knotted surfaces (see [4]), as well as in formulating set-theoretic solutions to the Yang-Baxter equation (see [5]). The binary operations of quandles generalize the notion of conjugation relations. In Wirtinger presentation, the fundamental group of a link is described using conjugation relations among the generators. There is an algebraic structure, called a conjugation quandle, where a group $G$ is equipped with the binary operation $x*y=y^{-1}xy$, forming a quandle on $G$.
This article will examine the residual finiteness of conjugation quandles. Before discussing the residual finiteness of quandles, we recall the concept of the residual finiteness. An algebraic structure $\mathcal{A}$ is called residually finite if for all $x,y\in\mathcal{A}$ with $x\neq y$, there is a finite algebraic structure $\mathcal{P}$ and a homomorphism $\psi:\mathcal{A}\rightarrow\mathcal{P}$ such that $\psi(x)\neq\psi(y)$. The residual finiteness of groups provides an important connection between group theory and topology. A group $G$ is residually finite if and only if the profinite topology on $G$ is Hausdorff. All finitely generated 3-manifold groups, including link groups, have been shown to be residually finite (see [10]). Similar to the residual finiteness of groups, these concepts can also be applied to quandles. It was first introduced in [2], where the authors proved that conjugation quandles are residually finite if their underlying groups are residually finite. Notably, the same authors subsequently showed that link quandles are also residually finite (see [3]). This is significant because it ensures that finite quandles can always be chosen for the coloring of links, which is a key tool for constructing link invariants.
Exploring the residual finiteness of conjugation quandles builds a bridge between geometric group theory and quandle theory. In geometric group theory, there is a challenging open problem posed by Gromov [9], concerning whether every hyperbolic group is residually finite. Since conjugation quandles are residually finite if their underlying groups are residually finite (see [2]), if there exists a non-residually finite conjugation quandle whose underlying group is hyperbolic, then there exists a non-residually finite hyperbolic group. Moreover, a common strategy is to use a less restrictive property to approach Gromov’s question, specifically by utilizing the Hopf property. An algebraic structure is called Hopfian if every surjective endomorphism of the structure is an automorphism. Mal’cev [12] proved that every finitely generated residually finite group is Hopfian. Since hyperbolic groups are finitely generated, researchers have examined Gromov’s question through the lens of the Hopf property. This perspective raised the question of whether there exist non-Hopfian hyperbolic groups. Recently, Reinfeldt and Weidmann [15] proved that all hyperbolic groups are Hopfian. Also subsequently, Fujiwara and Sela [8] provided an alternative proof, confirming this result. Although non-Hopfian hyperbolic groups do not exist, the Hopf property of quandles, which remains an area of limited exploration, still offers a potential approach to addressing Gromov’s question. In a way analogous to Mal’cev’s result, it is shown in [2] that every finitely generated residually finite quandle is Hopfian. Moreover, it is also established in [2] that conjugation quandles are residually finite if their underlying groups are residually finite, and the residual finiteness is subquandle-closed. Thus, we obtain the following.
  • If all hyperbolic groups are residually finite, then every finitely generated subquandle of a conjugation quandle with a hyperbolic underlying group is Hopfian.
In this article, we provide a complete characterization of the residual finiteness of conjugation quandles on the Baumslag-Solitar groups as stated in Theorem 3.1. To demonstrate this, we use Dehn quandles, which are finitely generated subquandles of conjugation quandles. The reason for using Dehn quandles lies in the major difference between groups and conjugation quandles derived from them. Specifically, many conjugation quandles are infinitely generated, even if their underlying groups are finitely generated. Regarding this, we show that the conjugation quandles of Baumslag-Solitar groups are always infinitely generated. In the latter part of the paper, as a result concerning the Hopf property of conjugation quandles, we showed that if a conjugation quandle is Hopfian, then its underlying group is also Hopfian. Accordingly, the conditions for non-Hopfian conjugation quandles of Baumslag-Solitar groups can be identified.
2. Preliminaries
In this section, we recall necessary definitions and known results of quandles. We begin with the definition of quandles.
Definition 2.1 (Quandles).
A non-empty set $Q$ equipped with a binary operation $*:Q\times Q\rightarrow Q$ is called a quandle if it satisfies the following three conditions:
  1. (i)
    $x*x=x$ for all $x\in Q$.
  2. (ii)
    For $x,y\in Q$, there exists a unique $z\in Q$ such that $x=z*y$.
  3. (iii)
    $(x*y)*z=(x*z)*(y*z)$ for all $x,y,z\in Q$.
The three conditions of a quandle $(Q,*)$ directly came from the Reidemeister moves of classical knots. In axiom (ii) of Definition 2.1, the element $z\in Q$ such that $x=z*y$ can be expressed by $x*^{-1}y$. Clearly, the inverse operation $*^{-1}$ also satisfies (i)-(iii), and thus the set $Q$ equipped with the binary operation $*^{-1}:Q\times Q\rightarrow Q$ is a quandle as well.
In the remainder of the article, we will use
$x_{1}*^{\epsilon_{1}}x_{2}*^{\epsilon_{2}}\cdots*^{\epsilon_{k-1}}x_{k}$
to denote
$(\cdots((x_{1}*^{\epsilon_{1}}x_{2})*^{\epsilon_{2}}x_{3})*\cdots)*^{\epsilon_{k-1}}x_{k}$
in a quandle, where $\epsilon_{i}=\pm 1$ for each $i=1,\dots,k-1$.
Quandles can be derived from groups. Especially, every group can be associated with a conjugation quandle as follows.
Definition 2.2 (Conjugation quandles).
Given a group $G$, the set $G$ equipped with the binary operation $x*y=y^{-1}xy$ which gives a quandle structure on $G$ is called the conjugation quandle of $G$, and is denoted by $\mathrm{Conj}(G)$.
Note that even though the underlying groups of conjugation quandles are finitely generated, this does not necessarily mean that the conjugation quandles themselves are finitely generated. We will discuss this further in Section 3. On the other hand, there exist finitely generated quandles that exhibit the same operation as conjugation quandles, as detailed below.
Definition 2.3 (Dehn quandles).
Let $G$ be a group, $X$ a nonempty subset of $G$ and $X^{G}$ the set of all conjugates of elements of $X$ in $G$. We call Dehn quandle of $G$ with respect to $X$, denoted by $\mathcal{D}(X^{G})$, the set $X^{G}$ equipped with the binary operation $x*y=y^{-1}xy$ which gives a quandle structure on $\mathcal{D}(X^{G})$.
Clearly, for each $X\subseteq G$, Dehn quandle $\mathcal{D}(X^{G})$ is a subquandle of $\mathrm{Conj}(G)$. Also, every subquandle of $\mathrm{Conj}(G)$ is a Dehn quandle (see [6, Corollary 3.9]). Moreover, if a group $G$ is generated by $X$, then $\mathcal{D}(X^{G})$ is generated as a quandle by $X$ (see [6, Proposition 3.2]).
In this article, we will examine the conjugation quandles of Baumslag-Solitar groups. Baumslag and Solitar [1] examined various groups with the presentations
$\mathrm{BS}(m,n)=\langle a,b\,|\,b^{-1}a^{m}b=a^{n}\rangle,$
where $m,n\in\mathbb{Z}\setminus\{0\}$. One of the most famous examples of a non-Hopfian group is the Baumslag-Solitar group $\mathrm{BS}(2,3)=\langle a,b\,|\,b^{-1}a^{2}b=a^{3}\rangle$. Also we can summarize their findings as follows (with a correction contributed by Meskin [14]):
  • $\mathrm{BS}(m,n)$ is residually finite if and only if $|m|=1$ or $|n|=1$ or $|m|=|n|$.
  • $\mathrm{BS}(m,n)$ is Hopfian if and only if it is residually finite or $\pi(m)=\pi(n)$, where $\pi(m)$ and $\pi(n)$ denote the set of all primes dividing $m$ and $n$, respectively.
3. Conjugation quandles of Baumslag-Solitar groups
The main result of this paper is a complete classification of the residual finiteness for conjugation quandles of the Baumslag-Solitar groups. A detailed proof of the following theorem will be provided in the latter part of Section 3.
Theorem 3.1.
Let $\mathrm{BS}(m,n)$ be the Baumslag-Solitar group given by the presentation
$\mathrm{BS}(m,n)=\langle a,b\,|\,b^{-1}a^{m}b=a^{n}\rangle,$
where $m,n\in\mathbb{Z}\setminus\{0\}$. Then the conjugation quandle $\mathrm{Conj}(\mathrm{BS}(m,n))$ is residually finite if and only if $|m|=1$, $|n|=1$ or $|m|=|n|$.
Every finitely generated residually finite quandle is Hopfian (see [2, Theorem 5.7]). However, the fact that conjugation quandles are non-Hopfian does not necessarily imply that they are always non-residually finite. This is due to the fact that, although the underlying groups of conjugation quandles are finitely generated, it does not guarantee that the conjugation quandles themselves are finitely generated. In detail, we provide an example of infinitely generated conjugation quandles along with a concrete proof.
Theorem 3.2.
Let $\mathrm{BS}(m,n)$ be the Baumslag-Solitar group given by the presentation
$\mathrm{BS}(m,n)=\langle a,b\,|\,b^{-1}a^{m}b=a^{n}\rangle,$
where $m,n\in\mathbb{Z}\setminus\{0\}$. Then the conjugation quandle $\mathrm{Conj}(\mathrm{BS}(m,n))$ is infinitely generated.
Proof.
Suppose to the contrary that $\mathrm{Conj}(\mathrm{BS}(m,n))$ is finitely generated. Let $M$ denote the number of elements in the generating set of $\mathrm{Conj}(\mathrm{BS}(m,n))$. Assume, for the sake of contradiction, that $M$ is bounded. The proof will be divided into four cases based on the values of $m$ and $n$.
Case 1. $|m|=1$ or $|n|=1$.
Since switching $m$ and $n$ leads to the same reasoning, we consider the case where $|m|=1$ without loss of generality. Let $L$ be the minimum number of generators of $\mathrm{Conj}(\mathrm{BS}(m,n))$ needed to generate both $a$ and $b$. Clearly, $L\leq M$. For every +$\ell\in\mathbb{Z}_{+}\cup\{0\}$, since $a^{\pm n^{\ell}}=a^{\pm 1}*b\underbrace{*\cdots*}_{\ell\text{-times}}b$, $a^{\pm n^{\ell}}$ can be generated using only the elements $a$ and $b$.
We use induction on $k$ to show that if $a^{q_{1}},\dots,a^{q_{k}}\in\mathrm{BS}(m,n)$ are the elements such that for each $j=1,\dots,k$, $q_{j}\neq\pm n^{\ell}$ for any +$\ell\in\mathbb{Z}_{+}\cup\{0\}$ and $q_{j}\neq\pm q_{t}n^{\ell}$ for $j\neq t$ with $t=1,\dots,k$, then the minimum number of generators of $\mathrm{Conj}(\mathrm{BS}(m,n))$ needed to generate $a^{q_{1}},\dots,a^{q_{k}}$ is at least $L+k$. To prove that this is true for $k=1$, we will show that $a^{q_{1}}$ cannot be generated using only the elements $a^{n^{\ell}}$ and $b$ such that +$\ell\in\mathbb{Z}_{+}\cup\{0\}$. Assuming the contrary, we have
(1) $a^{q_{1}}=x_{1}*\cdots*x_{h},$
where $x_{i}\in\{a^{\pm n^{\ell}},b^{\pm 1}\,|\,\ell\in\mathbb{Z}_{+}\cup\{0\}\}$ for each $i=1,\dots,h$. Then since
$a^{\pm n^{\ell_{1}}}*a^{\pm n^{\ell_{2}}}=a^{\pm n^{\ell_{1}}},\quad a^{\pm n^{\ell}}*b=a^{\pm n^{\ell+1}}\quad\text{and}\quad a^{\pm n^{\ell+1}}*^{-1}b=a^{\pm n^{\ell}},$
for every $\ell,\ell_{1},\ell_{2}\in\mathbb{Z}_{+}\cup\{0\}$, by replacing the left-hand sides $a^{\pm n^{\ell_{1}}}*a^{\pm n^{\ell_{2}}}$, $a^{\pm n^{\ell}}*b$ and $a^{\pm n^{\ell+1}}*^{-1}b$ with the right-hand sides $a^{\pm n^{\ell_{1}}}$, $a^{\pm n^{\ell+1}}$ and $a^{\pm n^{\ell}}$ respectively, we may assume $h$ to be the smallest number that satisfies equation (1). Hence there exists $r\in\mathbb{Z}_{+}$ such that
$a^{q_{1}}=g_{r}^{-1}b^{-\epsilon_{r}}g_{r-1}^{-1}b^{-\epsilon_{r-1}}\cdots g_{1}^{-1}b^{-\epsilon_{1}}g_{0}b^{\epsilon_{1}}g_{1}\cdots b^{\epsilon_{r-1}}g_{r-1}b^{\epsilon_{r}}g_{r}$
in $\mathrm{BS}(m,n)$, where $g_{i}\in\{1,a^{\pm n^{\ell}}\,|\,\ell\in\mathbb{Z}_{+}\cup\{0\}\}$ for every $i=0,1,\dots,r$. Since $g_{r}a^{q_{1}}g_{r}^{-1}=a^{q_{1}}$ in $\mathrm{BS}(m,n)$, we then have
(2) $a^{-q_{1}}b^{-\epsilon_{r}}g_{r-1}^{-1}b^{-\epsilon_{r-1}}\cdots g_{1}^{-1}b^{-\epsilon_{1}}g_{0}b^{\epsilon_{1}}g_{1}\cdots b^{\epsilon_{r-1}}g_{r-1}b^{\epsilon_{r}}=1$
in $\mathrm{BS}(m,n)$. Since $h$ is the smallest number that satisfies equation (1), the left-hand side of equation (2) is cyclically reduced. But then, $r=0$ by Britton’s Lemma. Therefore, $a^{q_{1}}=g_{0}$ in $\mathrm{BS}(m,n)$, which is a contradiction to the fact that $q_{1}\neq\pm n^{\ell}$ with +$\ell\in\mathbb{Z}_{+}\cup\{0\}$. Thus, $a^{q_{1}}$ cannot be generated using only the elements $a^{n^{\ell}}$ and $b$ such that +$\ell\in\mathbb{Z}_{+}\cup\{0\}$, and hence $a^{q_{1}}$ must be added to the generating set, requiring at least $L+1$ generators.
Now, suppose that this is true up to $k$, and that $q_{k+1}\neq\pm q_{j}n^{\ell}$ for every $j=1,\dots,k$ and $q_{k+1}\neq\pm n^{\ell}$.
Similarly, assume that
(3) $a^{q_{k+1}}=y_{1}*\cdots*y_{p},$
where $y_{i}\in\{a^{\pm q_{j}n^{\ell}},a^{\pm n^{\ell}},b^{\pm 1}\,|\,j\in\{1,\dots,k\},\ \ell\in\mathbb{Z}_{+}\cup\{0\}\}$ for each $i=1,\dots,p$. Then, by the analogous argument, we may assume that $p$ is the smallest number satisfying equation (3). Hence, there exists $s\in\mathbb{Z}_{+}$ such that
$a^{q_{k+1}}=h_{s}^{-1}b^{-\delta_{s}}h_{s-1}^{-1}b^{-\delta_{s-1}}\cdots h_{1}^{-1}b^{-\delta_{1}}h_{0}b^{\delta_{1}}h_{1}\cdots b^{\delta_{s-1}}h_{s-1}b^{\delta_{s}}h_{s}$
in $\mathrm{BS}(m,n)$, where $h_{i}\in\{1,a^{\pm q_{j}n^{\ell}},a^{\pm n^{\ell}}\,|\,j\in\{1,\dots,k\},\ \ell\in\mathbb{Z}_{+}\cup\{0\}\}$ for every $i=0,1,\dots,s$. Since $h_{s}a^{q_{k+1}}h_{s}^{-1}=a^{q_{k+1}}$ in $\mathrm{BS}(m,n)$,
(4) $a^{-q_{k+1}}b^{-\delta_{s}}h_{s-1}^{-1}b^{-\delta_{s-1}}\cdots h_{1}^{-1}b^{-\delta_{1}}h_{0}b^{\delta_{1}}h_{1}\cdots b^{\delta_{s-1}}h_{s-1}b^{\delta_{s}}=1$
in $\mathrm{BS}(m,n)$. Since $p$ is the smallest number that satisfies (3), the left-hand side of (4) is cyclically reduced. Moreover, by Britton’s Lemma, $s=0$. Thus, $a^{q_{k+1}}=h_{0}$ in $\mathrm{BS}(m,n)$, which contradicts to the fact that $q_{k+1}\neq\pm q_{j}n^{\ell}$ for every $j=1,\dots,k$ and $q_{k+1}\neq\pm n^{\ell}$. Therefore, $a^{q_{k+1}}$ cannot be generated using only the elements in $\{a^{q_{j}n^{\ell}},a^{n^{\ell}},b\,|\,j\in\{1,\dots,k\},\ \ell\in\mathbb{Z}_{+}\cup\{0\}\}$, and thus, $a^{q_{k+1}}$ must be added to the generating set, requiring at least $(L+k)+1$ generators. Since $k$ can be taken as arbitrarily large, the value $(L+k)+1$ is unbounded. Therefore, in Case 1, $M$ is unbounded, which implies that $\mathrm{Conj}(\mathrm{BS}(m,n))$ is infinitely generated.
Case 2. $|m|=|n|$.
In this case, $b^{-1}a^{m}b=a^{\pm m}$ in $\mathrm{BS}(m,n)$. Let $L$ be the minimum number of generators of $\mathrm{Conj}(\mathrm{BS}(m,n))$ needed to generate both $a^{m}$ and $b$. We use induction on $k$ to show that if $a^{q_{1}},\dots,a^{q_{k}}\in\mathrm{BS}(m,n)$ are the elements such that $q_{j}\neq\pm m$ and $q_{j}$ is distinct for each $j=1,\dots k$, then the minimum number of generators of $\mathrm{Conj}(\mathrm{BS}(m,n))$ needed to generate $a^{q_{1}},\dots,a^{q_{k}}$ is at least $L+k$. This holds for $k=1$ due to reasons analogous to the argument that $a^{q_{1}}$ in Case 1 cannot be generated solely by $a^{\pm n^{\ell}}$ and $b$, $a^{q_{1}}$ cannot be generated solely by $a^{m}$ and $b$.
Now, suppose this holds for all values up to $k$ and suppose that $q_{k+1}\neq\pm q_{j}$ for every $j=1,\dots,k$ and $q_{k+1}\neq\pm m$. Then, for reasons similar to the argument that $a^{q_{k+1}}$ must be added to the generating set in Case 1, $a^{q_{k+1}}$ in this case must also be added to the generating set. Therefore, at least $(L+k)+1$ generators are required. Since $k$ can be arbitrarily large, the value $(L+k)+1$ is unbounded. Consequently, in Case 2, $M$ is unbounded, which implies that $\mathrm{Conj}(\mathrm{BS}(m,n))$ has infinitely many generators.
Case 3. $1\neq|m|\neq|n|\neq 1$, and either $m\mid n$ or $n\mid m$.
Since switching $m$ and $n$ leads to the same argument, we may assume without loss of generality that $m\mid n$. Then $n=mr$ for some $r\in\mathbb{Z}\setminus\{0,\pm 1\}$. Let $L$ be the minimum number of generators of $\mathrm{Conj}(\mathrm{BS}(m,n))$ needed to generate $a^{m\cdot r^{\ell}}$ and $b$ such that +$\ell\in\mathbb{Z}_{+}\cup\{0\}$. We use induction on $k$ to show that if $a^{q_{1}},\dots,a^{q_{k}}\in\mathrm{BS}(m,n)$ are the elements such that each $q_{j}$ is a prime number, distinct, and satisfies $\gcd(q_{j},m)=1=\gcd(q_{j},n)$ for each $j=1,\dots,k$, then the minimum number of generators of $\mathrm{Conj}(\mathrm{BS}(m,n))$ needed to generate $a^{q_{1}},\dots,a^{q_{k}}$ is at least $L+k$. To show this holds for $k=1$, assuming the contrary, we have
(5) $a^{q_{1}}=x_{1}*\cdots*x_{h},$
where $x_{i}\in\{a^{\pm m\cdot r^{\ell}},b\,|\,\ell\in\mathbb{Z}_{+}\cup\{0\}\}$ for each $i=1,\dots,h$. Then
$a^{\pm m\cdot r^{\ell_{1}}}*a^{\pm m\cdot r^{\ell_{2}}}=a^{\pm m\cdot r^{\ell_{1}}},$
$a^{\pm m}*b\underbrace{*\cdots*}_{\ell\text{-times}}b=a^{\pm m\cdot r^{\ell}},$ $\quad a^{\pm m\cdot r^{\ell}}*^{-1}b\underbrace{*^{-1}\cdots*^{-1}}_{\ell\text{-times}}b=a^{\pm m},$
for every $\ell_{1},\ell_{2},\ell\in\mathbb{Z}_{+}\cup\{0\}$. In both cases above, as in Case 1, by replacing to right-hand sides, we may assume $h$ to be the smallest number that satisfies equation (5). Thus, due to reasons analogous to the argument that $a^{q_{1}}$ in Case 1 cannot be generated solely by $a^{\pm n^{\ell}}$ and $b$, $a^{q_{1}}$ cannot be generated solely by $a^{m\cdot r^{\ell}}$ and $b$. Therefore, $a^{q_{1}}$ must be added to the generating set, requiring at least $L+1$ generators. Now, suppose that this holds up to $k$, and assume that $q_{k+1}$ is a prime number such that $\gcd(q_{k+1},q_{j})=1$ for each $j=1,\dots,k$ and $\gcd(q_{k+1},m)=1=\gcd(q_{k+1},n)$. For reasons analogous to the argument that $a^{q_{k+1}}$ must be included in the generating set in Case 1, it follows that $a^{q_{k+1}}$ must also be included in the generating set in this case. Thus, at least $(L+k)+1$ generators are needed. Since there are infinitely many prime numbers, $k$ can grow arbitrarily large. Therefore, the value $(L+k)+1$ is unbounded. This implies that $M$ is unbounded, and thus, $\mathrm{Conj}(\mathrm{BS}(m,n))$ has infinitely many generators in Case 3.
Case 4. $1\neq|m|\neq|n|\neq 1$, and neither $m\mid n$ nor $n\mid m$.
In this case, let $L$ be the minimum number of generators of $\mathrm{Conj}(\mathrm{BS}(m,n))$ needed to generate $a^{m}$, $a^{n}$ and $b$. We use induction on $k$ to show that for $a^{q_{1}},\dots,a^{q_{k}}\in\mathrm{BS}(m,n)$, where each $q_{j}$ is a distinct prime satisfying $\gcd(q_{j},m)=1=\gcd(q_{j},n)$ for $j=1,\dots,k$, the minimum number of generators of $\mathrm{Conj}(\mathrm{BS}(m,n))$ required to generate $a^{q_{1}},\dots,a^{q_{k}}$ is at least $L+k$, as in Case 3. To demonstrate that this holds for $k=1$, let us assume the contrary, which leads to the following equation:
(6) $a^{q_{1}}=x_{1}*\cdots*x_{h},$
where $x_{i}\in\{a^{\pm m},a^{\pm n},b\}$. Then
$a^{\pm m}*a^{\pm m}=a^{\pm m}*a^{\pm n}=a^{\pm m},$ $\quad a^{\pm n}*a^{\pm m}=a^{\pm n}*a^{\pm n}=a^{\pm n},$
$a^{\pm m}*b=a^{\pm n},$ $\quad a^{\pm n}*b^{-1}=a^{\pm m}.$
By replacing to right-hand sides, we may assume $h$ to be the smallest number that satisfies equation (6). Thus, for reasons similar to the argument in Case 1, where $a^{q_{1}}$ cannot be generated solely by $a^{\pm n^{\ell}}$ and $b$, it follows that $a^{q_{1}}$ cannot be generated solely by $a^{m},a^{n}$ and $b$. Next, assume that the statement holds up to $k$, and let $q_{k+1}$ be a prime number such that $\gcd(q_{k+1},q_{j})=1$ for each $j=1,\dots,k$ and $\gcd(q_{k+1},m)=1=\gcd(q_{k+1},n)$. Then, for reasons similar to those in Case 3, the value $(L+k)+1$ is unbounded. This implies that $M$ is unbounded, and hence, $\mathrm{Conj}(\mathrm{BS}(m,n))$ has infinitely many generators in Case 4.
After considering all four cases, this completes the proof of Theorem 3.2. ∎
Next, we proceed to state the following theorem concerning the Hopf property of conjugation quandles and their underlying groups.
Theorem 3.3.
Let $G$ be a group. If $\mathrm{Conj}(G)$ is Hopfian, then the group $G$ is Hopfian.
Proof.
Suppose to the contrary that $G$ is non-Hopfian. Then there exists an epimorphism $\psi:G\rightarrow G$ such that $\psi$ is not injective, which implies that $\psi$ induces the quandle epimorphism $\tilde{\psi}:\mathrm{Conj}(G)\rightarrow\mathrm{Conj}(G)$ which is not injective. Hence $\mathrm{Conj}(G)$ is non-Hopfian, which is a contradiction. Thus, $G$ is Hopfian. ∎
Using Theorem 3.3, given a non-Hopfian group, one can construct a non-Hopfian conjugation quandle whose underlying group is the given group. In particular, by Theorem 3.3, we directly obtain the following fact.
Corollary 3.4.
Let $\mathrm{BS}(m,n)$ be the Baumslag-Solitar group given by the presentation
$\mathrm{BS}(m,n)=\langle a,b\,|\,b^{-1}a^{m}b=a^{n}\rangle$
with $|m|\neq 1\neq|n|$ and $\pi(m)\neq\pi(n)$, where $\pi(m)$ and $\pi(n)$ denote the set of all primes dividing $m$ and $n$, respectively. Then the conjugation quandle $\mathrm{Conj}(\mathrm{BS}(m,n))$ is non-Hopfian.
Proof.
Since $\mathrm{BS}(m,n)$ is non-Hopfian if $|m|\neq 1\neq|n|$ and $\pi(m)\neq\pi(n)$, where $\pi(m)$ and $\pi(n)$ denote the set of all primes dividing $m$ and $n$, respectively, then by Theorem 3.3, it follows that $\mathrm{Conj}(\mathrm{BS}(m,n))$ is non-Hopfian. ∎
Now we give the proof of Theorem 3.1. In the proof of Theorem 3.1, we show that the conjugation quandle $\mathrm{Conj}(\mathrm{BS}(m,n))$ is non-residually finite if $1\neq|m|\neq|n|\neq 1$, by proving that $\mathrm{Conj}(\mathrm{BS}(m,n))$ contains a finitely generated subquandle which is non-Hopfian.
Proof of Theorem 3.1.
Suppose that $|m|=1$, $|n|=1$ or $|m|=|n|$. Then, since $\mathrm{BS}(m,n)$ with $|m|=1$, $|n|=1$ or $|m|=|n|$ is residually finite, by [2, Proposition 4.1], $\mathrm{Conj}(\mathrm{BS}(m,n))$ is residually finite.
Now, suppose that $1\neq|m|\neq|n|\neq 1$. We categorize this assumption into three distinct cases. In each case, we will derive that $\mathrm{Conj}(\mathrm{BS}(m,n))$ is non-residually finite. In this proof, let $\pi(m)$ and $\pi(n)$ denote the set of all primes dividing $m$ and $n$, respectively.
Case 1. $\pi(m)\neq\pi(n)$.
Let $\mathcal{D}$ be the Dehn quandle of $\mathrm{BS}(m,n)$ with respect to $\{a,b,a^{m}\}$. We will show that $\mathcal{D}$ is non-residually finite, which implies that $\mathrm{Conj}(\mathrm{BS}(m,n))$ is also non-residually finite due to the fact that the residual finiteness is subquandle-closed. To be specific, we will prove that $\mathcal{D}$ is finitely generated and non-Hopfian, thereby proving that $\mathcal{D}$ is non-residually finite. We will first establish the following claim.
Claim. The quandle $\mathcal{D}$ is finitely generated.
Proof of Claim.
Let $w\in\mathcal{D}$ be an arbitrary element. Then $w=g^{-1}xg$ for some $x\in\{a,b,a^{m}\}$ and $g\in\mathrm{BS}(m,n)$. If $g=1$ in $\mathrm{BS}(m,n)$, then clearly $w=x\in\{a,b,a^{m}\}$. Assume that $g\neq 1$ in $\mathrm{BS}(m,n)$. Since $g\in\mathrm{BS}(m,n)$, we may write $g\equiv x_{1}^{\epsilon_{1}}\cdots x_{k}^{\epsilon_{k}}$ such that $x_{i}\in\{a,b\}$ and $\epsilon_{i}=\pm 1$ for each $i=1,\dots,k$. Hence $w$ can be written as $x_{k}^{-\epsilon_{k}}\cdots x_{1}^{-\epsilon_{1}}x\;x_{1}^{\epsilon_{1}}\cdots x_{k}^{\epsilon_{k}}$, where $x\in\{a,b,a^{m}\}$, $x_{i}\in\{a,b\}$, and $\epsilon_{i}=\pm 1$ for each $i=1,\dots,k$. In other words, $w$ can be expressed as either
$a*^{\epsilon_{1}}x_{1}*^{\epsilon_{2}}\cdots*^{\epsilon_{k}}x_{k},\quad b*^{\epsilon_{1}}x_{1}*^{\epsilon_{2}}\cdots*^{\epsilon_{k}}x_{k},\quad\text{or}\quad a^{m}*^{\epsilon_{1}}x_{1}*^{\epsilon_{2}}\cdots*^{\epsilon_{k}}x_{k},$
where $x_{i}\in\{a,b\}$ and $\epsilon_{i}=\pm 1$ for each $i=1,\dots,k$. Thus, $\mathcal{D}$ is generated by $\{a,b,a^{m}\}$. ∎
Now we verify that $\mathcal{D}$ is non-Hopfian. Let $\phi$ be the unique quandle homomorphism from a free quandle on $\{a,b,a^{m}\}$ to $\mathcal{D}$ induced by the mapping
$a\mapsto a^{m},\quad b\mapsto b\quad\text{and}\quad a^{m}\mapsto a.$
Then $\phi$ induces an endomorphism $\tilde{\phi}$ of $\mathcal{D}$ and clearly, $\tilde{\phi}$ is surjective. However, $\tilde{\phi}$ is not injective since $\tilde{\phi}(a*b*a*^{-1}b)=\tilde{\phi}(a)$ but $a*b*a*^{-1}b\neq a$ in $\mathcal{D}$ since $ba^{-1}b^{-1}abab^{-1}\neq a$ in $\mathrm{BS}(m,n)$ by Britton’s Lemma. Therefore, $\mathcal{D}$ is non-residually finite, and thus, for Case 1, $\mathrm{Conj}(\mathrm{BS}(m,n))$ is non-residually finite.
Case 2. Either $m\mid n$ or $n\mid m$, and $\pi(m)=\pi(n)$.
Without loss of generality, we may assume that $n=\ell m$ for some $\ell\in\mathbb{Z}\setminus\{0,\pm 1\}$. Let $\mathcal{D}$ be the Dehn quandle of $\mathrm{BS}(m,n)$ with respect to $\{a,b,a^{m}\}$. We will demonstrate that $\mathcal{D}$ is non-residually finite by applying a similar argument to that used in Case 1. In analogy with the conclusion drawn in Case 1, it follows that $\mathcal{D}$ is finitely generated.
Now we prove that $\mathcal{D}$ is non-Hopfian. Let $\phi$ be the unique homomorphism from a free quandle on $\{a,b,a^{m}\}$ to $\mathcal{D}$ induced by the mapping
$a\mapsto a,\quad b\mapsto b\quad\text{and}\quad a^{m}\mapsto a^{m}*b.$
So $\phi$ induces an endomorphism $\tilde{\phi}$ of $\mathcal{D}$. Clearly, $\tilde{\phi}$ is surjective. However $\tilde{\phi}$ is not injective since
$\tilde{\phi}(a^{m}*^{-1}b*a)=\tilde{\phi}(a^{m}*^{-1}b)$
but $a^{m}*^{-1}b*a\neq a^{m}*^{-1}b$ in $\mathcal{D}$, since $a^{-1}ba^{m}b^{-1}a\neq ba^{-m}b^{-1}$ in $\mathrm{BS}(m,n)$ by Britton’s Lemma. Thus, for Case 2, $\mathrm{Conj}(\mathrm{BS}(m,n))$ is non-residually finite.
Case 3. Neither $m\mid n$ nor $n\mid m$, while $\pi(m)=\pi(n)$.
In this case, since $\pi(m)=\pi(n)$, there exists $k=\gcd(m,n)$ with $k\neq 1$. Then
$\pi\left(\frac{m}{k}\right)\neq\pi\left(\frac{n}{k}\right)\quad\text{and}\quad\left|\frac{m}{k}\right|\neq 1\neq\left|\frac{n}{k}\right|.$
Hence, by the conclusion of Case 1, $\mathrm{Conj}(\mathrm{BS}(\frac{m}{k},\frac{n}{k}))$ is non-residually finite. Since $\mathrm{BS}(\frac{m}{k},\frac{n}{k})$ is a subgroup of $\mathrm{BS}(m,n)$, and since the residual finiteness is subgroup-closed, it follows that $\mathrm{Conj}(\mathrm{BS}(m,n))$ is non-residually finite for Case 3. This end the proof of Theorem 3.1. ∎
By Theorem 3.2, the conjugation quandle $\mathrm{Conj}(\mathrm{BS}(m,n))$ is always infinitely generated. Thus, it is not possible to observe the Hopf property using Theorem 3.1 based on the fact that finitely generated residually finite quandles are Hopfian (see [2]). In this regard, we leave the following question.
Question 3.5.
If $m,n\in\mathbb{Z}\setminus\{0\}$ with $|m|=1$, $|n|=1$, or $|m|=|n|$, is the conjugation quandle $\mathrm{Conj}(\mathrm{BS}(m,n))$ Hopfian?
Finally, in order to achieve a complete classification of both the residual finiteness and the Hopf property, we pose the following question.
Question 3.6.
If $m,n\in\mathbb{Z}\setminus\{0\}$ with $\pi(m)=\pi(n)$, where $\pi(m)$ and $\pi(n)$ denote the set of all primes dividing $m$ and $n$, respectively, then is the conjugation quandle $\mathrm{Conj}(\mathrm{BS}(m,n))$ Hopfian?
Acknowledgement
Mohamed Elhamdadi was partially supported by Simons Foundation collaboration grant 712462. Jan Kim was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (Grant No. 2019R1A6A1A11051177).